# Segmented Operations using Matrix Multiplications

### v1

@@ 4.2 Putting everything together
Algorithm 1 speculatively computes the “local” unsegmented scans of each block of size $s$. Then, the speculated scan will be updated through a $\mathsf{REVSPEC}$ vector instruction, which corrects the misspeculated values to complete the local segmented scans.
@@
9:    $\bm{\mathrm{f}}_{s}\leftarrow\mathsf{NOT}\left(\mathsf{ISZERO}\left(\overline{\bm{\mathrm{f}}}_{s}\right)\right)$triangleright View flags in $\{0,1\}$
@@
17:    $\overline{\bm{\mathrm{x}}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{x}},\bm{\mathrm{U}}_{s})$triangleright Speculative $s$-block scan
@@
19:    return $\mathsf{REVSPEC}\left(\overline{\bm{\mathrm{x}}},\overline{\bm{\mathrm{f}}};s\right)$triangleright Revert speculation to form segmented scan
@@
23:    $\bm{\mathrm{m}}\leftarrow\mathsf{ISZERO}\left(\mathsf{SUB}\left(\overline{\bm{\mathrm{f}}},\overline{\bm{\mathrm{f}}}_{start}\right)\right)$triangleright $\bm{\mathrm{m}}(i)=1$ if $\overline{\bm{\mathrm{f}}}(i)$ is equal to the first element of its segment
24:    $\bm{\mathrm{m}}\leftarrow\mathsf{AND}\left(\bm{\mathrm{m}},\neg\bm{\mathrm{f}}\right)$triangleright No update on first element of each segment
25:    $\bm{\mathrm{w}}\leftarrow\mathsf{FILLS}\left(\bm{\mathrm{z}};s\right)$triangleright $\bm{\mathrm{w}}(is:(i+1)s)=\bm{\mathrm{z}}(is)$ (unfiltered correction vector)
26:    $\bm{\mathrm{w}}\leftarrow\mathsf{MASK}\left(\bm{\mathrm{w}},\bm{\mathrm{m}}\right)$triangleright Filtered (masked) correction vector
@@
Algorithm 1 Segmented scan in the MMV-RAM model
In particular, at each recursive step, the algorithm computes the “local” segmented scan of each $s$-block of the input $\bm{\mathrm{x}}$ in two phases (BlockSegScan in Algorithm 1). In the first phase, we compute the unsegmented scans of each block by speculative matrix multiplications using $\bm{\mathrm{U}}_{s}$ as the right matrix operand, see Line 17 of Algorithm 1. In the second phase, the algorithm reverts any misspeculated segment using vector operations ($\mathsf{REVSPEC}$ in Algorithm 1). Given an unsegmented scan of any $s$-block, Lemma 4.1 describes an AC[0] circuit for $\mathsf{REVSPEC}$, which reverts the misspeculated scans. We highlight that such a circuit is only achievable because we have already precomputed the unsegmented scans of $\bm{\mathrm{x}}$ and $\bm{\mathrm{f}}$.
Lemma 4.1.
Proof.
The proof is detailed in Section A.2. ∎
To use Lemma 4.1 in Algorithm 1, recall that the speculation prepares the entire vectors $\overline{\bm{\mathrm{x}}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{x}},\bm{\mathrm{U}}_{s})$ and $\overline{\bm{\mathrm{f}}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{f}},\bm{\mathrm{U}}_{s})$, which contain $\mathcal{O}(n/s)$ blocks of size $s$. By concatenating $\lceil n/s\rceil$ copies of the circuit of Lemma 4.1, we can build a uniform family of circuits that implements the desired vector instruction $\mathsf{REVSPEC}$. The total circuit size is $\mathcal{O}\left(\frac{n}{s}(s^{2}B+B^{2})\right)=\mathcal{O}\left(n(sB+B^{2}/s)\right)$, and the depth remains constant after concatenation.
In the last phase of Algorithm 1, following the recursive call, the VCU are invoked to scalar-broadcast-vector-add the previous block’s last entry (scalar) to the current block’s first segment (UpdateFirstSegment in Algorithm 1).
Figure 4.1 depicts an example of an execution of Algorithm 1 for an input of size $16$ and block size $s=4$. In this example, matrix multiplication computes the unsegmented scans, and the misspeculated entries are coloured red. The reversion of the speculation is followed before the recursions take place. In the final step, vector-masked operations are employed to propagate the last entry of each block to the first segment of the next block.
Remark 4.2.
The circuit of Lemma 4.1 is ideal for our theoretical analysis, but it is rather complicated compared to the rest of the vector instructions in Table 3.3. In a realistic setting, it would need to be implemented as a dedicated circuit inside the VCU, which can be prohibitive in terms of area. In Section A.3 we describe an alternative, Algorithm 2, which simulates the $\mathsf{REVSPEC}$ instruction using only a few standard vector instructions, which are typically implemented in existing architectures.
4.3 Analysis
Here, we analyze Algorithm 1 in terms of the number of steps (depth), the number of vector operations, and the number of matrix multiplications. First, we note that if the input flag vector $\bm{\mathrm{f}}$ is equal to zero, we recover the unsegmented scan algorithm (prefix sum) of [64], which is a generalization of the Brent–Kung algorithm [11]. In this case, we can completely avoid calling $\mathsf{REVSPEC}$. In Theorem 4.3 we report the complexity of the algorithm of [64] in the MMV-RAM model, which was originally analyzed in the TCU model. The work-depth bounds match the ones of [64] with respect to $n$ and $s$.

### v2

@@ A.2 Block-Recursive Segmented Scan Algorithm
Algorithm˜1 speculatively computes the “local” unsegmented scans of each block of size $s$. Then, the speculated scan will be updated through a $\mathsf{REVSPEC}$ vector instruction, which corrects the misspeculated values to complete the local segmented scans.
@@
9:  $\bm{\mathrm{f}}_{s}\leftarrow\mathsf{NOT}\left(\mathsf{ISZERO}\left(\overline{\bm{\mathrm{f}}}_{s}\right)\right)$ $\triangleright$ View flags in $\{0,1\}$
@@
17:  $\overline{\bm{\mathrm{x}}}\leftarrow\mathsf{MATMUL}(\bm{\mathrm{x}},\bm{\mathrm{U}}_{s})$ $\triangleright$ Speculative $s$-block scan
@@
19:  return $\mathsf{REVSPEC}\left(\overline{\bm{\mathrm{x}}},\overline{\bm{\mathrm{f}}};s\right)$ $\triangleright$ Revert speculation to form segmented scan
@@
23:  $\bm{\mathrm{m}}\leftarrow\mathsf{ISZERO}\left(\mathsf{SUB}\left(\overline{\bm{\mathrm{f}}},\overline{\bm{\mathrm{f}}}_{start}\right)\right)$ $\triangleright$ $\bm{\mathrm{m}}(i)=1$ if $\overline{\bm{\mathrm{f}}}(i)$ is equal to the first element of its segment
24:  $\bm{\mathrm{m}}\leftarrow\mathsf{AND}\left(\bm{\mathrm{m}},\neg\bm{\mathrm{f}}\right)$ $\triangleright$ No update on first element of each segment
25:  $\bm{\mathrm{w}}\leftarrow\mathsf{FILLS}\left(\bm{\mathrm{z}};s\right)$ $\triangleright$ $\bm{\mathrm{w}}(is:(i+1)s)=\bm{\mathrm{z}}(is)$ (unfiltered correction vector)
26:  $\bm{\mathrm{w}}\leftarrow\mathsf{MASK}\left(\bm{\mathrm{w}},\bm{\mathrm{m}}\right)$ $\triangleright$ Filtered (masked) correction vector
@@
Algorithm 1 Block-recursive segmented scan in the MMV-RAM model